From GlueXWiki
Jump to: navigation, search

April 2, 2015 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Running at half energy
  2. Gas puzzle
    • CO2 monitoring [2]
  3. CDC update (Beni, Mike)
  4. FDC update (Lubomir)
  5. Electronics (Fernando, Cody, Chris, Nick)
  6. Engineering
  7. Other


Participants: Curtis, Naomi(CMU), Sean(NU), Eugene, Luke, Mike, Dave, Chris, Nick, Cody, Beni, Simon, and Lubomir (JLab).

Running at half energy

- For the FDC we expect to have about the same e.m. background on the chambers, i.e. we will be limited by the same photon intensity as in the fall run. The number of pion tracks is expected to be about half of that in the fall. We asked for 16 hours beam time assuming we will have 8 hours at each of the two HV settings: 2200V and 2250V. As the HV filter boxes are expected to reduce the noise significantly, we ask for 16h run only at one setting - 2200V but at a lower discriminator threshold - thus compensating for the smaller number of tracks.

Gas puzzle

- The new pump has been installed in the EEL and data taken with the spare package using the certified gas mixture. On the attached plot: CO2 measurements (blue), atmospheric pressure (black), corrected measurements (red) using (dp/p)^2.4 pressure correction. This pump has variable flow, Nick and Chris made voltage control for that. Thus the flow was adjusted in this way, keeping the rotameter open completely, so that the pump doesn't create vacuum.

- The plan is to monitor the certified gas till Friday afternoon, then move the sensor to the Hall and monitor the CDC for one more day. On Saturday afternoon Beni will change the CO2 gas factor to 1.

CDC update

- Naomi found in run 2451, some of the pedestals are wobbling by 100+/-3 units on a scale of hours. Could be temperature variations, in any case it is small compared to the thresholds of ~60.

FDC update

- The HV channel on FDC4 cell 6 sector 2p, developed a constant current of ~3 microamps (at the nominal 2200V) after being unstable for a week (current/voltage plots attached). Most likely the problem is outside of the gas volume and we may attempt to fix it sometimes when the FDC is pulled outside of the magnet.

- Interesting results with an 55Fe source shown on the plots: ratio of the upper strip to down strip charge as a function of the avalanche position along the wire used in the trigger for both tests, with the spare package and with FDC4 in the Hall. Apart from variations along the wire caused by some differences in the gains of the strips and position w.r.t. center of the strip, there are two images 10% apart. Beni proposed these are avalanches above and below the wire, but the avalanches develop so close to the wire that is difficult to explain the 10% difference.


- The HV filter boxes will arrive next week and will be installed right away.


- We discussed if we can monitor spikes/glitches in the power, since we suspect some of the HV trips are related to that. Dave has such monitoring for the magnet, but needs to be put into the EPICS stream.